Atsie (atsie):

ry + s = tx - m for y. I have to explain each step in this solution. How would I do so?

OpenStudy (aajc10):

what grade are you in

Atsie (atsie):

Whale, thats a long story. I'm in 10th but I'm currently doing 9th grade math. A bit far behind.

OpenStudy (faiqraees):

Okay so you have to isolate y. Start by subtracting s on both sides

OpenStudy (faiqraees):

I cant really help, if you dont respond

OpenStudy (aajc10):

yep

Atsie (atsie):

@FaiqRaees Sorry about that, I walked away for a bit. So subtracting s from both sides would look like \[s - s + s - y\] Right? Or wrong?

OpenStudy (faiqraees):

ry + s -s= tx - m -s

Atsie (atsie):

Oh, okay. What would the next step be then?

Atsie (atsie):

@FaiqRaees :)

OpenStudy (faiqraees):

So now your equation is ry= tx - m -s right?

Atsie (atsie):

Correct. We have multiply something next?

OpenStudy (faiqraees):

Since r is being multiplied with y we have to divide r on both sides to get rid of r on right hand side

Atsie (atsie):

Divide r by what? @FaiqRaees

Atsie (atsie):

@MissSmartiez Could you finish helping me? I'd really appreciate it! ^.^

OpenStudy (misssmartiez):

Sorry hun, I am in a rush to get this stuff done, and I don't think I can explain. ^

Atsie (atsie):

Its okay, I understand. :)

Atsie (atsie):

@welshfella Could you help me at all? ^.^

OpenStudy (anonymous):

i can help, i found a website that has the answers to these kinds of problems

OpenStudy (anonymous):

like, no effort at all

OpenStudy (anonymous):

if u want it i can share

OpenStudy (texaschic101):

ry = tm - x - s...divide both sides by r

Atsie (atsie):

@charlotte321 I'd rather at least try and learn how to work through a math problem first before I head on chase an answer. Thank you though anyways. :) @texaschic101 Both sides meaning...the r and the t ?

OpenStudy (texaschic101):

yes...the r and the t side...basically, both sides of the = sign

OpenStudy (faiqraees):

ry/r= (tx - m -s)/r

rebeccaxhawaii (rebeccaxhawaii):

get y by its self

Atsie (atsie):

@texaschic101 Alright, and then what would be the end result? Like as far as solving for the next part of it?

OpenStudy (texaschic101):

end result is y = (tm - x - s) / r

OpenStudy (texaschic101):

because the r's on the left side cancel

Atsie (atsie):

@texaschic101 Your right, the r's do cancel so that would make sense. Thank you so much!! ^.^

OpenStudy (texaschic101):

yw